8 gram of oxygen gas at S.T.P. is expanded so that the volume is doubled. the work done during this process is.

Grade:11

1 Answers

Gopal
6 years ago
number of moles of 8g oxygen = 0.25 moles
during epansion of gas  volume is doubled V2= 2V1
Work done= -2.303 . nRTlog\frac{V2}{V1}
Work done= -2.303 (0.25)RTlog\frac{2V1}{V1}
W=-0.5757RT log2
=-0.173RT joul